OpenStudy (anonymous):

what is the area of a triangle if the base is 4 1/2 ft and the height is 2 2/3 ft? what is the area of a triangle if the base is 1.2 m and the height is 0.6 m?

OpenStudy (jdoe0001):

\(\bf \textit{Area of a Triangle}=\cfrac{{\color{blue}{ base}}\times {\color{green}{ height}}}{2} \\ \quad \\ \quad \\ \cfrac{{\color{blue}{ 4\frac{1}{2}}}\times {\color{green}{ 2\frac{2}{3}}}}{2}\qquad and \qquad \cfrac{{\color{blue}{ 1.2}}\times {\color{green}{ 0.6}}}{2}\)